# Revisiting Graph Analytics Benchmark

### v1

@@ 3.1. Algorithm Popularity
Table 2 presents statistics on the number of papers related to the selected algorithms, appearing in representative conferences and journals (Meng et al., 2024). It also shows their frequency of appearance in research engines like DBLP, Google Scholar, and Web of Science (WoS) over the past ten years. Moreover, the selected algorithms are not only frequently studied in academic papers but also have a wide range of practical applications. Here are just a few examples: (1) Social Network Analysis. $\mathsf{PR}$ and $\mathsf{BC}$ are used to detect important individuals in social networks by considering the influence of neighbors (Jin et al., 2012; Yin et al., 2019; Gao et al., 2022) and shortest paths (Kourtellis et al., 2013), respectively. $\mathsf{LPA}$ and $\mathsf{KC}$ are used to detect communities (Azaouzi and Romdhane, 2017; Ouyang et al., 2020), while $\mathsf{CD}$ can further reveal the hierarchical structure (Malliaros et al., 2020). (2) Computer Vision. $\mathsf{WCC}$ can find connected pixels for unsupervised segmentation (Wang, 1998; Trevor et al., 2013). $\mathsf{LPA}$ can utilize given labels for semantic segmentation (Vernaza and Chandraker, 2017; Mustikovela et al., 2016). (3) Biological Analysis. Similar to the social network analysis, $\mathsf{BC}$ and $\mathsf{PR}$ are also adaptive to measure the importance of proteins (Wang et al., 2024; Iván and Grolmusz, 2011), while $\mathsf{TC}$ can measure the similarity like Jaccard (Buriol et al., 2006). In epidemiology, $\mathsf{CD}$ is suitable to reveal and predict epidemic outbreaks. (4) Road Network Routing. On road network, $\mathsf{SSSP}$ and $\mathsf{WCC}$ are used to compute shortest paths (Forster and Nanongkai, 2018; Nanongkai, 2014) and check the connection (Xie and Levinson, 2007; Wiedemann and Ebner, 2000) between vertices, respectively. $\mathsf{BC}$ is used to detect frequent area that most shortest paths passed (LI et al., 2010; Kazerani and Winter, 2009). (5) Recommendation Systems. $\mathsf{PR}$ can detect high influence items for recommendation (Ma et al., 2008). $\mathsf{TC}$ and $\mathsf{KC}$ can detect clique and triadic instance to mine a close relation for further recommendation (Wu et al., 2016; Carullo et al., 2015; Caramia and Felici, 2006). $\mathsf{KC}$ and $\mathsf{LPA}$ can also detect fraud information (Berry et al., 2004; Li et al., 2022).
@@ 3.2. Algorithm Diversity
Table 3 further compares the computational workload and topics of our core algorithms and $\mathsf{LDBC}$. Our algorithms span a range of complexities, from linear ($\mathsf{PR}$, $\mathsf{CD}$, $\mathsf{WCC}$, $\mathsf{LPA}$) to logarithmic-linear ($\mathsf{SSSP}$), polynomial ($\mathsf{TC}$, $\mathsf{BC}$), and even higher-order complexities ($\mathsf{KC}$), providing a balanced coverage of different workload levels. In contrast, more than half of the algorithms ($\mathsf{PR}$, $\mathsf{LPA}$, $\mathsf{WCC}$, $\mathsf{BFS}$) in $\mathsf{LDBC}$ are linear. Regarding the topics, our core algorithms cover areas including centrality, community detection, traversal, cohesive subgraph, and pattern matching, highlighting the diversity of our algorithms, while $\mathsf{LDBC}$ only includes centrality, community detection, traversal, with half of its algorithms focused on community detection. Overall, our core algorithms are not only widely used in academia and industry, reflecting real-world needs, but also cover diverse computation workloads and topics, ensuring our benchmark can effectively identify platform bottlenecks and strengths.
@@ 3.3. Algorithm Suitability for Computing Models
Graph Computing Models. The Bulk Synchronous Parallelism (BSP) model, introduced by Valiant (Valiant, 1990), is a popular computing model for parallel graph computation. It divides the computation into multiple supersteps, where each superstep consists of three phases: local computation, communication, and synchronization, for efficiently executing parallel tasks.
@@
On top of the BSP model, the $\mathsf{vertex}$-$\mathsf{centric}$ model was developed with the “Think-Like-a-Vertex” (TLAV) philosophy  (McCune et al., 2015). In $\mathsf{vertex}$-$\mathsf{centric}$ model, the vertex is the basic unit of computing and scheduling. Each vertex can process computation locally and send messages to other vertices. Users are only required to implement a vertex-based computation interface, while the platform will execute it over all vertices for a certain number of iterations or until it converges. There are a few variants of $\mathsf{vertex}$-$\mathsf{centric}$ model such as Single-Phase ($\mathsf{Pregel}$ (Malewicz et al., 2010), $\mathsf{Pregel+}$ (Yan et al., 2015), $\mathsf{Flash}$ (Li et al., 2023b), $\mathsf{Ligra}$ (Shun and Blelloch, 2013b)), $\mathsf{Scatter}$-$\mathsf{Gather}$ (Signal/Collect (Stutz et al., 2010)), and $\mathsf{Scatter}$-$\mathsf{Combine}$ (GRE (Yan et al., 2014b)). Specially, Spark also supports $\mathsf{vertex}$-$\mathsf{centric}$ by providing Pregel-like interfaces, i.e., $\mathsf{VertexRDD}$ in GraphX (Gonzalez et al., 2014). $\mathsf{Vertex}$-$\mathsf{centric}$ models face potential challenges, including load imbalance, which becomes particularly problematic in power-law graphs, and significant communication overhead when deployed on large-scale clusters.
@@
Some platforms build upon the $\mathsf{vertex}$-$\mathsf{centric}$ model by extending computation on edges or groups of vertices, leading to the development of $\mathsf{edge}$-$\mathsf{centric}$ and $\mathsf{block}$-$\mathsf{centric}$ models, respectively. PowerGraph (Gonzalez et al., 2012), $\mathsf{X}$-$\mathsf{Stream}$ (Roy et al., 2013), $\mathsf{Graphchi}$ (Kyrola et al., 2012) and $\mathsf{Chaos}$ (Roy et al., 2015) provide an $\mathsf{edge}$-$\mathsf{centric}$ model to execute tasks over edges to resolve load skew in power-law graphs and fully utilize sequential reading and writing of solid-slate disk. Alternatively, $\mathsf{Blogel}$ (Yan et al., 2014a) and $\mathsf{Grape}$ (Fan et al., 2017a) are $\mathsf{block}$-$\mathsf{centric}$ platforms that divide the graph into multiple blocks, such that vertex functions can communicate without communication if they belong to the same block. However, $\mathsf{edge}$-$\mathsf{centric}$ does not support communication with a non-neighbor, while it’s often non-trivial to program with the $\mathsf{block}$-$\mathsf{centric}$ model.
@@
All the above computing models produce output sizes proportional to the graph size, making them unsuitable for graph-mining problems that may yield an exponential number of results (Yan et al., 2017, 2024). To address this, platforms like $\mathsf{Arabesque}$ (Teixeira et al., 2015), $\mathsf{Fractal}$ (Dias et al., 2019), $\mathsf{AutoMine}$ (Mawhirter and Wu, 2019), $\mathsf{Peregrine}$ (Jamshidi et al., 2020), and $\mathsf{G}$-$\mathsf{thinker}$ (Yan et al., 2020) adopt the $\mathsf{subgraph}$-$\mathsf{centric}$ model. In this model, the fundamental computing unit is the subgraph (e.g., a triangle) rather than vertices, vertex groups, or edges. Users need to define how to construct candidate subgraphs and implement a computation interface for each subgraph.
@@
Algorithm Categorizations. Different graph computing models can greatly influence algorithm efficiency, as their diverse algorithmic requirements may align differently with the characteristics of these models. Drawing on insights from prior studies (Li et al., 2023a; Yan et al., 2014a), we categorize the eight selected algorithms into three main classes and analyze their varying suitability across different computing models.
@@
  • Sequential Algorithms ($\mathsf{SSSP}$, $\mathsf{WCC}$, $\mathsf{BC}$, $\mathsf{CD}$) require a specific execution order and often involve frequent synchronizations, such as iterations in $\mathsf{SSSP}$. The $\mathsf{block}$-$\mathsf{centric}$ model is particularly well-suited for such algorithms, especially those with fewer iterations. While it is possible to implement algorithms using the $\mathsf{vertex}$-$\mathsf{centric}$ and $\mathsf{edge}$-$\mathsf{centric}$ models, this can lead to suboptimal performance. However, some $\mathsf{vertex}$-$\mathsf{centric}$-based platforms (Li et al., 2023b; Shun and Blelloch, 2013b) can execute computation on a vertex subset instead of activate all vertices to minimize unnecessary computations.
@@ 4. Data Generator & Synthetic Datasets
Among various data generators, the LDBC Graphalytics generator ($\mathsf{LDBC}$-$\mathsf{DG}$) is noted for its high-quality simulation of real-life social graphs (Prat-Pérez and Dominguez-Sal, 2014). Specifically, $\mathsf{LDBC}$-$\mathsf{DG}$ (1) initially generates vertices with some properties (such as location and interest). (2) Then, vertices are sorted with some specific similarity metrics (e.g., Z-order for location and identifier order for interest). (3) Finally, edges are formed with probabilities influenced by the distances between vertices in this ordered sequence (Erling et al., 2015).
Briefly, $\mathsf{LDBC}$-$\mathsf{DG}$ follow the “Homophily Principle” (McPherson et al., 2001) that individuals with similar characteristics are more likely to form connections. Therefore, similar vertices are closer in distance and have a higher probability of being connected. $\mathsf{LDBC}$-$\mathsf{DG}$ defines the edge formation probability between vertex $u_{i}$ and $u_{j}$ as follows: $\Pr[e(u_{i},u_{j})]=max\{p^{|j-i|},p_{limit}\}$ , where $p$ is the base probability, $i$ and $j$ represents the position of $u_{i}$ and $u_{j}$ in the ordered sequence, $p_{limit}$ is a lower-bound. By default, $p$ and $p_{limit}$ are $0.95$ and $0.2$, respectively. To generate edges, $\mathsf{LDBC}$-$\mathsf{DG}$ iterates over each vertex $u_{i}$ and progressively samples $u_{j}(j>i)$ with the probability $\Pr[e(u_{i},u_{j})]$ as shown in Figure 1 until the number of required edges is obtained. However, $\mathsf{LDBC}$-$\mathsf{DG}$ has two key limitations:
  • Inflexibility: $\mathsf{LDBC}$-$\mathsf{DG}$ focuses on producing social networks of different scales. However, graphs of the same scale can differ in characteristics like density and diameter, affecting platform performance. For instance, subgraph algorithms are sensitive to graph density, while sequential algorithms can be impacted by diameter. Some platforms offer optimizations for high-density (Yan et al., 2015; Fan et al., 2018) or large-diameter graphs (Yan et al., 2014a). To effectively evaluate platform performance, the data generator should provide the flexibility of creating graphs with varied characteristics.
@@ 4.1. Failure-Free Trial Generator
As shown in Figure 2, we introduce a new probability function. For a vertex $u_{i}$, the probability of connecting an edge to $u_{j}$ $(j>i)$ is defined as $\frac{1}{c+(j-i)}$. $c\geq 0$ is an adjustable parameter, where a larger $c$ can reduce the probability of each edge. In $\mathsf{FFT}$-$\mathsf{DG}$, $c$ defaults to 00 such that the adjacent edge $e(u_{i},u_{i+1})$ always exists. As verified in Section 8.1, this new function does not affect the quality of generated graphs compared to that of $\mathsf{LDBC}$-$\mathsf{DG}$.
@@
(1) $\Pr[e(i,j)]$ $=\left(1-\frac{1}{c+1}\right)\cdot\left(1-\frac{1}{c+2}\right)...\frac{1}{c+(j-i)}$
@@
Then, the probabilities for each edge to be the first existing edge are structured such that the sequence of probabilities, $1-\frac{c}{c+1}$, $\frac{c}{c+1}-\frac{c}{c+2}$, etc., forms a continuous range that spans from 00 to $1$. By treating these probabilities as distinct segments that collectively cover the interval $(0,1]$, we can randomly select a floating-point number $f$ from this interval $(0,1]$ to represent an edge $e(i,j)$ by $j=i+\left\lfloor\left(\frac{1}{f}-1\right)\cdot c\right\rfloor+1$. Once an edge is sampled (e.g., $e(u_{i},u_{j})$), the following edges $e(u_{i},u_{k})\ (k{>}j)$ exhibit the similar properties:
@@
The probabilities for each edge to become the next existing edge are structured similarly as $1-\frac{c^{\prime}}{c^{\prime}+1}$, $\frac{c^{\prime}}{c^{\prime}+1}-\frac{c^{\prime}}{c^{\prime}+2}$, and so forth. Consequently, we can consistently obtain connected edges by updating the parameter $c$ without any failure samples. The details are outlined in Algorithm 1.

### v2

@@ 3.1. Algorithm Popularity
Table 2 presents statistics on the number of papers related to the selected algorithms, appearing in representative conferences and journals (meng2024survey). It also shows their frequency of appearance in research engines like DBLP, Google Scholar, and Web of Science (WoS) over the past ten years. Moreover, the selected algorithms are not only frequently studied in academic papers but also have a wide range of practical applications. Here are just a few examples: (1) Social Network Analysis. $\mathsf{PR}$ and $\mathsf{BC}$ are used to detect important individuals in social networks by considering the influence of neighbors (jin2012lbsnrank; yin2019signed; gao2022mebc) and shortest paths (kourtellis2013identifying), respectively. $\mathsf{LPA}$ and $\mathsf{KC}$ are used to detect communities (azaouzi2017evidential; ouyang2020clique), while $\mathsf{CD}$ can further reveal the hierarchical structure (malliaros2020core). (2) Computer Vision. $\mathsf{WCC}$ can find connected pixels for unsupervised segmentation (wang1998stochastic; trevor2013efficient). $\mathsf{LPA}$ can utilize given labels for semantic segmentation (vernaza2017learning; mustikovela2016can). (3) Biological Analysis. Similar to the social network analysis, $\mathsf{BC}$ and $\mathsf{PR}$ are also adaptive to measure the importance of proteins (wang2024efficient; ivan2011web), while $\mathsf{TC}$ can measure similarity metrics like Jaccard similarity (buriol2006counting). In epidemiology, $\mathsf{CD}$ is suitable to reveal and predict epidemic outbreaks. (4) Road Network Routing. In road networks, $\mathsf{SSSP}$ and $\mathsf{WCC}$ are used to compute shortest paths (forster2018faster; nanongkai2014distributed) and check the connection (xie2007measuring; wiedemann2000automatic) between vertices, respectively. $\mathsf{BC}$ is applied to identify high-traffic areas frequently traversed by shortest paths (li2010betweenness; kazerani2009can). (5) Recommendation Systems. $\mathsf{PR}$ can detect high influence items for recommendation (ma2008bringing). $\mathsf{TC}$ and $\mathsf{KC}$ can detect cliques and triadic patterns to uncover close relationships for recommendation (wu2016link; carullo2015triadic; caramia2006mining). $\mathsf{KC}$ and $\mathsf{LPA}$ are also used for fraud detection (berry2004emergent; li2022adaptive).
@@ 3.2. Algorithm Diversity
Table 3 further compares the computational workload and topics of our core algorithms and $\mathsf{LDBC}$. Our algorithms span a range of complexities, from linear ($\mathsf{PR}$, $\mathsf{CD}$, $\mathsf{WCC}$, $\mathsf{LPA}$) to logarithmic-linear ($\mathsf{SSSP}$), polynomial ($\mathsf{TC}$, $\mathsf{BC}$), and even higher-order complexities ($\mathsf{KC}$), providing a balanced coverage of different workload levels. In contrast, more than half of the algorithms ($\mathsf{PR}$, $\mathsf{LPA}$, $\mathsf{WCC}$, $\mathsf{BFS}$) in $\mathsf{LDBC}$ are linear. Regarding the topics, our core algorithms cover areas including centrality, community detection, traversal, cohesive subgraph, and pattern matching, highlighting the diversity of our algorithms, while $\mathsf{LDBC}$ only includes centrality, community detection, traversal, with half of its algorithms focused on community detection. Overall, our core algorithms are widely used in academia and industry, addressing real-world needs while covering diverse workloads, enabling our benchmark to identify platform bottlenecks and strengths effectively.
@@ 3.3. Algorithm Suitability for Computing Models
Graph Computing Models. The Bulk Synchronous Parallelism (BSP) model, introduced by Valiant (valiant1990bridging), is a popular computing model for parallel graph computation. It divides the computation into multiple supersteps, where each superstep consists of three phases: local computation, communication, and synchronization, for efficiently executing parallel tasks.
@@
On top of the BSP model, the $\mathsf{vertex}$-$\mathsf{centric}$ model was developed with the “Think-Like-a-Vertex” (TLAV) philosophy (mccune2015thinking). In $\mathsf{vertex}$-$\mathsf{centric}$ model, the vertex is the basic unit of computing and scheduling. Each vertex can process computation locally and send messages to other vertices. Users are only required to implement a vertex-based computation interface, while the platform will execute it over all vertices for a certain number of iterations or until it converges. There are a few variants of $\mathsf{vertex}$-$\mathsf{centric}$ model such as Single-Phase ($\mathsf{Pregel}$ (malewicz2010pregel), $\mathsf{Pregel+}$ (yan2015effective), $\mathsf{Flash}$ (10184838), $\mathsf{Ligra}$ (DBLP:conf/ppopp/ShunB13)), $\mathsf{Scatter}$-$\mathsf{Gather}$ (Signal/Collect (stutz2010signal)), and $\mathsf{Scatter}$-$\mathsf{Combine}$ (GRE (yan2014pregel)). Specially, Spark also supports $\mathsf{vertex}$-$\mathsf{centric}$ by providing Pregel-like interfaces, i.e., $\mathsf{VertexRDD}$ in GraphX (gonzalez2014graphx). $\mathsf{Vertex}$-$\mathsf{centric}$ models face potential challenges, including load imbalance, which becomes particularly problematic in power-law graphs, and significant communication overhead when deployed on large-scale clusters.
@@
Some platforms build upon the $\mathsf{vertex}$-$\mathsf{centric}$ model by extending computation on edges or groups of vertices, leading to the development of $\mathsf{edge}$-$\mathsf{centric}$ and $\mathsf{block}$-$\mathsf{centric}$ models, respectively. PowerGraph (gonzalez2012powergraph), $\mathsf{X}$-$\mathsf{Stream}$ (roy2013x), $\mathsf{Graphchi}$ (kyrola2012graphchi) and $\mathsf{Chaos}$ (roy2015chaos) provide an $\mathsf{edge}$-$\mathsf{centric}$ model to execute tasks over edges to resolve load skew in power-law graphs and fully utilize sequential reading and writing of solid-state disk. Alternatively, $\mathsf{Blogel}$ (yan2014blogel) and $\mathsf{Grape}$ (fan2017grape) are $\mathsf{block}$-$\mathsf{centric}$ platforms that divide the graph into multiple blocks, such that vertex functions within the same block can exchange information without communication. However, $\mathsf{edge}$-$\mathsf{centric}$ models do not support non-neighbor communication, while it’s often non-trivial to program with the $\mathsf{block}$-$\mathsf{centric}$ model.
@@
All the above computing models produce output sizes proportional to the graph size, making them unsuitable for graph-mining problems that may yield an exponential number of results (yan2017big; yan2024systems). To address this, platforms like $\mathsf{Arabesque}$ (teixeira2015arabesque), $\mathsf{Fractal}$ (dias2019fractal), $\mathsf{AutoMine}$ (mawhirter2019automine), $\mathsf{Peregrine}$ (jamshidi2020peregrine), and $\mathsf{G}$-$\mathsf{thinker}$ (yan2020g) adopt the $\mathsf{subgraph}$-$\mathsf{centric}$ model. In this model, the fundamental computing unit is the subgraph (e.g., a triangle) rather than vertices, vertex groups, or edges. Users need to define how to construct candidate subgraphs and implement a computation interface for each subgraph.
@@
Algorithm Categorizations. Different graph computing models can greatly influence algorithm efficiency, as their diverse algorithmic requirements may align differently with the characteristics of these models. Drawing on insights from prior studies (li2023flash; yan2014blogel), we categorize the eight selected algorithms into three main classes and analyze their suitability across different computing models.
@@
  • Sequential Algorithms ($\mathsf{SSSP}$, $\mathsf{WCC}$, $\mathsf{BC}$, $\mathsf{CD}$) require a specific execution order and often involve frequent synchronizations, such as iterations in $\mathsf{SSSP}$. The $\mathsf{block}$-$\mathsf{centric}$ model is particularly well-suited for such algorithms, especially those with fewer iterations. While it is possible to implement algorithms using the $\mathsf{vertex}$-$\mathsf{centric}$ and $\mathsf{edge}$-$\mathsf{centric}$ models, this can lead to suboptimal performance. However, some $\mathsf{vertex}$-$\mathsf{centric}$-based platforms (10184838; DBLP:conf/ppopp/ShunB13) can execute computation on a vertex subset instead of activating all vertices to minimize unnecessary computations.
@@ 4. Data Generator & Synthetic Datasets
Among various data generators, the LDBC Graphalytics generator ($\mathsf{LDBC}$-$\mathsf{DG}$) is noted for its high-quality simulation of real-life social graphs (prat2014community). Specifically, $\mathsf{LDBC}$-$\mathsf{DG}$ (1) initially generates vertices with some properties (such as location and interest). (2) Then, vertices are sorted with some specific similarity metrics (e.g., Z-order for location and identifier order for interest). (3) Finally, edges are formed with probabilities influenced by the distances between vertices in this ordered sequence (erling2015ldbc).
Briefly, $\mathsf{LDBC}$-$\mathsf{DG}$ follows the “Homophily Principle” (mcpherson2001birds) that individuals with similar characteristics are more likely to form connections. Therefore, similar vertices are closer in distance and have a higher probability of being connected. $\mathsf{LDBC}$-$\mathsf{DG}$ defines the edge formation probability between vertex $u_{i}$ and $u_{j}$ as follows: $\Pr[e(u_{i},u_{j})]=max\{p^{|j-i|},p_{limit}\}$ , where $p$ is the base probability, $i$ and $j$ represent the position of $u_{i}$ and $u_{j}$ in the ordered sequence, $p_{limit}$ is a lower-bound. By default, $p$ and $p_{limit}$ are $0.95$ and $0.2$, respectively. To generate edges, $\mathsf{LDBC}$-$\mathsf{DG}$ iterates over each vertex $u_{i}$ and progressively samples $u_{j}(j>i)$ with the probability $\Pr[e(u_{i},u_{j})]$ as shown in Figure 1 until the number of required edges is obtained. However, $\mathsf{LDBC}$-$\mathsf{DG}$ has two key limitations:
  • Inflexibility: $\mathsf{LDBC}$-$\mathsf{DG}$ focuses on producing social networks of different scales. However, graphs of the same scale can differ in characteristics like density and diameter, affecting platform performance. For instance, subgraph algorithms are sensitive to graph density, while sequential algorithms can be impacted by diameter. Some platforms offer optimizations for high-density (yan2015effective; fan2018parallelizing) or large-diameter graphs (yan2014blogel). To effectively evaluate platform performance, the data generator should provide the flexibility of creating graphs with varied characteristics.
@@ 4.1. Failure-Free Trial Generator
As shown in Figure 2, we introduce a new probability function. For a vertex $u_{i}$, the probability of connecting an edge to $u_{j}$ $(j>i)$ is defined as $\frac{1}{c+(j-i)}$. $c\geq 0$ is an adjustable parameter, where a larger $c$ can reduce the probability of each edge. In $\mathsf{FFT}$-$\mathsf{DG}$, $c$ defaults to 0 such that the adjacent edge $e(u_{i},u_{i+1})$ always exists. As verified in Section 8.1, this new function does not affect the quality of generated graphs compared to that of $\mathsf{LDBC}$-$\mathsf{DG}$.
@@
(1) $\Pr[e(i,j)]$ $=\left(1-\frac{1}{c+1}\right)\cdot\left(1-\frac{1}{c+2}\right).\frac{1}{c+(j-i)}$
@@
Then, the probabilities for each edge to be the first existing edge are structured such that the sequence of probabilities, $1-\frac{c}{c+1}$, $\frac{c}{c+1}-\frac{c}{c+2}$, etc., forms a continuous range that spans from 0 to $1$. By treating these probabilities as distinct segments that collectively cover the interval $(0,1]$, we can randomly select a floating-point number $f$ from this interval $(0,1]$ to represent an edge $e(i,j)$ by $j=i+\left\lfloor\left(\frac{1}{f}-1\right)\cdot c\right\rfloor+1$. Once an edge is sampled (e.g., $e(u_{i},u_{j})$), the following edges $e(u_{i},u_{k})\ (k{>}j)$ exhibit similar properties:
@@
The probabilities for each edge to become the next existing edge are structured similarly to $1-\frac{c^{\prime}}{c^{\prime}+1}$, $\frac{c^{\prime}}{c^{\prime}+1}-\frac{c^{\prime}}{c^{\prime}+2}$, and so forth. Consequently, we can consistently obtain connected edges by updating the parameter $c$ without any failure samples. The details are outlined in Algorithm 1.